# The Dynamic, the Static, and the Weak factor models and the analysis of high-dimensional time series

### v1

@@ 8 Conclusions
(5) To conclude, let us show that (ii) ${\bm{\Gamma}}^{(n)}_{\chi}$ has exactly $r$ divergent eigenvalues while (iii) the eigenvalues of ${\bm{\Gamma}}^{(n)}_{\xi}$ all are bounded. Let us assume that the number of diverging eigenvalues of ${\bm{\Gamma}}^{(n)}_{\chi}$ is strictly less than $r$. Since ${\bm{\Gamma}}^{(n)}_{X}$ decomposes into ${\bm{\Gamma}}^{(n)}_{\chi}+{\bm{\Gamma}}^{(n)}_{\xi}$ and has $r$ diverging eigenvalues, the Weyl inequalities imply that the first eigenvalue $\lambda^{(n)}_{\xi;1}$ of ${\bm{\Gamma}}^{(n)}_{\xi}$ tends to infinity as $n\to\infty$. Then, there exists a sequence $(w^{(n)}_{\xi})_{n\in{\mathbb{N}}}$ with $w^{(n)}_{\xi}\coloneqq\sum_{i=1}^{n}c^{(n)}_{i}\xi_{it}\in{\mathcal{H}}_{\scriptscriptstyle{\text{\rm idio}}}$ and $\sum_{i=1}^{n}(c^{(n)}_{i})^{2}=1$ such that the variance $(\sigma^{(n)}_{\xi})^{2}$ of $w^{(n)}_{\xi}$ tends to infinity as $n\to\infty$. Since $\xi_{it}$ is the projection of $X_{it}$ on ${\mathcal{H}}_{\scriptscriptstyle{\text{\rm idio}}}$$w^{(n)}_{\xi}$ also is a linear combination $\sum_{i=1}^{n}d^{(n)}_{i}X_{it}$ of $X_{1t},\ldots,X_{nt}$ and (the eigenvalues of a nondegenerate projection operator are either 1 or zero, and cannot all be zero) $0<C^{-}\leq\sum_{i=1}^{n}(d^{(n)}_{i})^{2}\leq 1$. Hence, the sequence $(w^{(n)}_{\xi*}\coloneqq w^{(n)}_{\xi}/\sigma^{(n)}_{\xi})_{n\in{\mathbb{N}}}$, where each $w^{(n)}_{\xi*}$ has variance one, is in ${\mathcal{W}}_{*}$. This bounded sequence contains a subsequence converging weakly to some $w_{\xi*}$; since each $w^{(n)}_{\xi*}$ is in ${\mathcal{H}}_{\scriptscriptstyle{\text{\rm idio}}}$, hence orthogonal to ${\mathcal{H}}_{\scriptscriptstyle{\text{\rm com}}}$, so is the weak limit $w_{\xi*}$. Proceeding as in (3), it can be shown that this $w_{\xi*}$ also is the strong limit of a sequence of elements of ${\mathcal{H}}_{r\rceil}$, hence belongs to ${\mathcal{H}}_{\scriptscriptstyle{\text{\rm com}}}$; moreover, since the convergence is strong, Var($w_{\xi*}$)=1. The intersection of ${\mathcal{H}}_{\scriptscriptstyle{\text{\rm com}}}$ and ${\mathcal{H}}_{\scriptscriptstyle{\text{\rm idio}}}$ being $\{0\}$, this implies $w_{\xi*}=0$. This, however, is incompatible with Var($w_{\xi*}$)=1. Therefore, the number of diverging eigenvalues of ${\bm{\Gamma}}^{(n)}_{\chi}$ cannot be strictly less than $r$. The Weyl inequalities and the fact that ${\bm{\Gamma}}^{(n)}_{X}={\bm{\Gamma}}^{(n)}_{\chi}+{\bm{\Gamma}}^{(n)}_{\xi}$ then imply that all eigenvalues of ${\bm{\Gamma}}^{(n)}_{\xi}$ are bounded as $n\to\infty$. Claims (ii) and (iii) follow.

### v3

@@ 10 Conclusions
(5) To conclude, let us show that (ii) ${\bm{\Gamma}}^{(n)}_{\chi}$ has exactly $r$ divergent eigenvalues while (iii) the eigenvalues of ${\bm{\Gamma}}^{(n)}_{\xi}$ all are bounded. Let us assume that the number of diverging eigenvalues of ${\bm{\Gamma}}^{(n)}_{\chi}$ is strictly less than $r$. Since ${\bm{\Gamma}}^{(n)}_{X}$ decomposes into ${\bm{\Gamma}}^{(n)}_{\chi}+{\bm{\Gamma}}^{(n)}_{\xi}$ and has $r$ diverging eigenvalues, the Weyl inequalities imply that the first eigenvalue $\lambda^{(n)}_{\xi;1}$ of ${\bm{\Gamma}}^{(n)}_{\xi}$ tends to infinity as $n\to\infty$. Then, there exists a sequence $(w^{(n)}_{\xi})_{n\in{\mathbb{N}}}$ with $w^{(n)}_{\xi}\coloneqq\sum_{i=1}^{n}c^{(n)}_{i}\xi_{it}\in{\mathcal{H}}_{\scriptscriptstyle{\text{\rm idio}}}$ and $\sum_{i=1}^{n}(c^{(n)}_{i})^{2}=1$ such that the variance $(\sigma^{(n)}_{\xi})^{2}$ of $w^{(n)}_{\xi}$ tends to infinity as $n\to\infty$. Since $\xi_{it}$ is the projection of $X_{it}$ on ${\mathcal{H}}_{\scriptscriptstyle{\text{\rm idio}}}$$w^{(n)}_{\xi}$ also is a linear combination $\sum_{i=1}^{n}d^{(n)}_{i}X_{it}$ of $X_{1t},\ldots,X_{nt}$ and (the eigenvalues of a nondegenerate projection operator are either 1 or zero, and cannot all be zero) $0<C^{-}\leq\sum_{i=1}^{n}(d^{(n)}_{i})^{2}\leq 1$. Hence, the sequence $(w^{(n)}_{\xi*}\coloneqq w^{(n)}_{\xi}/\sigma^{(n)}_{\xi})_{n\in{\mathbb{N}}}$, where each $w^{(n)}_{\xi*}$ has variance one, is in ${\mathcal{W}}_{*}$. This bounded sequence contains a subsequence converging weakly to some $w_{\xi*}$; since each $w^{(n)}_{\xi*}$ is in ${\mathcal{H}}_{\scriptscriptstyle{\text{\rm idio}}}$, hence orthogonal to ${\mathcal{H}}_{\scriptscriptstyle{\text{\rm com}}}$, so is the weak limit $w_{\xi*}$. Proceeding as in (3), it can be shown that this $w_{\xi*}$ also is the strong limit of a sequence of elements of ${\mathcal{H}}_{r\rceil}$, hence belongs to ${\mathcal{H}}_{\scriptscriptstyle{\text{\rm com}}}$; moreover, since the convergence is strong, Var($w_{\xi*}$)=1. The intersection of ${\mathcal{H}}_{\scriptscriptstyle{\text{\rm com}}}$ and ${\mathcal{H}}_{\scriptscriptstyle{\text{\rm idio}}}$ being $\{0\}$, this implies $w_{\xi*}=0$. This, however, is incompatible with Var($w_{\xi*}$)=1. Therefore, the number of diverging eigenvalues of ${\bm{\Gamma}}^{(n)}_{\chi}$ cannot be strictly less than $r$. The Weyl inequalities and the fact that ${\bm{\Gamma}}^{(n)}_{X}={\bm{\Gamma}}^{(n)}_{\chi}+{\bm{\Gamma}}^{(n)}_{\xi}$ then imply that all eigenvalues of ${\bm{\Gamma}}^{(n)}_{\xi}$ are bounded as $n\to\infty$. Claims (ii) and (iii) follow. $\square$